OpenStudy (anonymous):

What volume of 0.100 M Ca(OH)2 is required to neutralize 100.0 mL of a 0.200 M HCI solution?

OpenStudy (anonymous):

I do, but I am creating a quizlet AP chemistry study guide that has over 400 questions and answers. I dont feel like doing the work for all of them so I aquire help on here. I would apreciate an asnwer but if not I will just work it out myself.

OpenStudy (anonymous):

First, write and balance the equation Second, work out the molar ratio of Ca(OH)2 to HCl. According to the question, HCl is the limiting reagent. The equation will be Ca(OH)2 + 2HCl - CaCl2 + 2H2O The molar ratio is 1:2 To find the moles of the acid, we use the formula n=cv/1000, where n is mol, c is concentration and v is volume in cm^3. n= 0.2*100/1000=0.02mol of HCl Then the mol of the base will be 0.02*1/2= 0.01 mol Using the mol formula , making v the subject will be v=1000*n/c= 1000*0.01/0.1=100mL.

OpenStudy (anonymous):

Haha thanks a lot mate but I already worked this one out :P

OpenStudy (anonymous):

No problem.